# Masked Autoencoders that Feel the Heart: Unveiling Simplicity Bias for ECG Analyses

### v1

@@ Abstract
The diagnostic value of electrocardiogram (ECG) lies in its dynamic characteristics, ranging from rhythm fluctuations to subtle waveform deformations that evolve across time and frequency domains. However, supervised ECG models tend to overfit dominant and repetitive patterns, overlooking fine-grained but clinically critical cues—a phenomenon known as Simplicity Bias (SB), where models favor easily learnable signals over subtle but informative ones. In this work, we first empirically demonstrate the presence of SB in ECG analyses and its negative impact on diagnostic performance, while simultaneously discovering that self-supervised learning (SSL) can alleviate it, providing a promising direction for tackling the bias. Following the SSL paradigm, we propose a novel method comprising two key components: 1) Temporal-Frequency aware Filters to capture temporal-frequency features reflecting the dynamic characteristics of ECG signals, and 2) building on this, Multi-Grained Prototype Reconstruction for coarse and fine representation learning across dual domains, further mitigating SB. To advance SSL in ECG analyses, we curate a large-scale multi-site ECG dataset with 1.53 million recordings from over 300 clinical centers. Experiments on three downstream tasks across six ECG datasets demonstrate that our method effectively reduces SB and achieves state-of-the-art performance. Code and dataset will be released publicly.
@@ 1 Introduction
The electrocardiogram (ECG) remains a cornerstone of non-invasive cardiac assessment, offering a cost-effective and widely accessible method for monitoring life-threatening conditions such as acute myocardial ischemia [1, 2, 3, 4]. The beauty of ECG lies in its structured patterns encoded across both temporal and frequency domains. As illustrated in Fig. 1, ECG signals exhibit prominent, stable patterns like QRS complexes and rhythm regularities, characterized as low-frequency features representing dominant coarse-grained characteristics. In contrast, rapidly varying and delicate patterns, such as P and T waves—crucial for detecting early-stage or complex cardiac abnormalities—are identified as high-frequency features, encoding subtle fine-grained diagnostic information. Despite the success of powerful supervised models in advancing ECG analyses [5, 6, 7, 8, 9], they often overfit to low-frequency patterns while underrepresenting high-frequency cues critical for clinical interpretation, a phenomenon known as simplicity bias [10, 11] in supervised learning paradigm.
Simplicity bias (SB) refers to the tendency of supervised neural networks to prioritize dominant, easily captured features while overlooking subtle but clinically important deviations [10, 12]. Conversely, self-supervised learning (SSL) can mitigate SB by encouraging models to learn more comprehensive representations [12, 13]. To investigate whether SB manifests in ECG analyses and whether existing ECG SSL methods can alleviate it, we conduct preliminary experiments. As illustrated in Fig. 2(a), supervised models such as ResNet [9] and ViT [14] tend to focus primarily on dominant QRS complexes and global pseudo-periodic rhythms, while often neglecting subtle high-frequency components such as P and fibrillatory (F) waves. In contrast, the self-supervised model (MAE [15]) demonstrates partially improved attention to these clinically significant regions, as further discussed in Section 2. Nonetheless, current ECG SSL methods [14, 15, 16, 17, 3, 18, 19] are not explicitly designed to address SB. In particular, they lack mechanisms to capture multi-granular representations and model the hierarchical dependencies across temporal-frequency domains, limiting their ability to fully capture the diagnostic richness embedded in ECG signals.
@@
To address such limitation, we propose a novel self-supervised learning method, muLti-grained tEmporal-frequency Aware maSked auToencoders (LEAST). Our proposed method, LEAST, integrates temporal-frequency attention and multi-grained modeling into a masked autoencoder framework, encouraging the self-supervised model to capture challenging features and mitigate SB. The Method consists of two key modules: (1) Temporal-Frequency aware Filters (TFF): We propose multi-head filters within the encoder to capture inherent patterns across both temporal and frequency domains. This design seamlessly integrates temporal and frequency information while capturing domain-specific patterns, enhancing the model’s sensitivity to the dynamic nature of ECG signals and effectively mitigating SB. Building on this, (2) Multi-Grained Prototype Reconstruction (MGPR): To further extend the model’s sensitivity from coarse-grained to fine-grained features, we construct prototypical representations at multiple semantic granularities across dual domains, encouraging the model to preserve information beyond dominant patterns. To advance self-supervised learning research in ECG analyses, we introduce a Chinese multi-site ECG arrhythmia dataset with approximately 380,000 samples collected from 356 clinical centers. Additionally, we integrate multiple publicly available datasets [20, 21] to build a large-scale multi-site ECG dataset, comprising a total of 1,530,000 samples.
@@ 2.1 Preliminary Experiments
We perform standard supervised learning on CPSC2018 [22] for ECG classification, using ResNet [9] as the backbone. Guided by class activation mapping (CAM) [23], we visualize the model’s attention on an atrial fibrillation (AFiB) signal (lead II used) in Fig. 2(a). The results reveal that the supervised model predominantly focuses on easily recognizable QRS complexes, while largely ignoring subtle regions like P waves and F waves, providing clear evidence of SB in ECG analyses.
@@ 2.2 Incorporating Self-Supervised Learning to Mitigate Simplicity Bias
Based on the previous verification of SB in supervised ECG models, we further evaluate whether SSL can alleviate this issue. We adopt a SSL framework in MAE fashion [15] and compare it with a supervised baseline, ViT [14], under identical settings (see Appendix A.1). As shown in Fig. 2(a), both models focus on dominant QRS complexes. However, the MAE shows moderately enhanced attention to the presence of P waves, partially mitigating SB. Nonetheless, the improvement remains limited, the ECG SSL model still underattends fine-grained structures critical for diagnosing complex arrhythmias.
@@ 4.2 Complementary Public ECG Datasets
MIMIC-IV-ECG [20] includes approximately 800,000 12-lead recordings from 160,000 patients collected between 2008–2019 at the Beth Israel Deaconess Medical Center in Boston, USA. As a representative of high-acuity inpatient and ICU care in a developed country, it offers valuable insight into ECG patterns under intensive monitoring.
CODE-15% [21] contains 345,779 12-lead ECGs from 233,770 patients across hundreds of municipalities in Brazil. This dataset reflects low-resource outpatient and primary care settings, contributing substantial geographic and epidemiological diversity.
@@ 5.1 Datasets and Empirical Settings
To advance research on self-supervised pretraining in ECG analyses, we constructed large-scale multi-site ECG pretraining dataset. It includes 380,000 proprietary samples collected from over 356 hospitals, fully annotated with 61 diagnostic categories, as well as curated public datasets [20, 21], resulting in a combined corpus of 1,530,000 samples. For fair comparison, all self-supervised baselines [15, 24, 25, 26, 17, 19] are re-pretrained on this dataset in our experiments.
For downstream adaptation, we adopt two paradigms: linear probing and full fine-tuning. We conduct experiments across three downstream tasks—classification, segmentation (detecting characteristic waveform components, i.e. R-peaks locating task) and forecasting (predicting survival time from ECG signals, a form of survival analyses)—on six widely used ECG datasets. More details about the task definition, downstream datasets and metrics could be seen in Appendix B.1. For a fair comparison, following [14, 15, 19], all experiments are conducted using the ViT-Base encoder [14], pretrained on our curated ECG dataset. The empirical settings for training are detailed in the Appendix B.4.
@@ 5.2 Main Results
We conduct classification experiments on PTB-XL [7], Ningbo [27], Chapman [28], and CPSC2018 [22], covering a wide range of clinical conditions, geographic regions, and patient demographics. We report the results in Table 2. Results show that our LEAST method consistently achieves significant accuracy improvements over state-of-the-art methods.
@@
MAE [25] 87.64 91.37 70.48 73.16 89.25 76.54 73.31 89.36 75.35 94.81 95.68 76.11
MOCO v3 [26] 86.76 90.92 67.17 73.40 90.12 76.11 74.41 86.13 74.93 92.37 93.70 70.42
MTAE [15] 86.79 89.46 66.41 72.89 88.49 76.63 75.33 89.44 76.24 95.15 95.46 73.00
MLAE [15] 86.68 88.87 63.88 71.61 88.35 74.35 69.69 88.26 71.12 94.16 94.90 65.67
PatchTST [24] 85.37 86.27 64.54 72.20 87.18 73.41 67.92 85.03 68.87 94.33 95.00 70.38
ST-MEM [19] 87.33 90.30 66.39 73.25 89.22 76.03 73.12 89.91 75.44 95.24 95.42 75.17
@@
MAE [25] 88.81 93.00 73.36 86.01 96.64 86.70 88.20 97.33 89.61 96.60 97.23 84.50
MOCO v3 [26] 87.31 92.27 71.45 86.55 96.42 84.39 85.25 96.61 87.12 95.28 95.29 79.24
MTAE [15] 87.39 90.33 69.36 86.31 95.52 87.45 85.94 97.70 85.06 95.37 95.73 83.04
MLAE [15] 88.12 90.28 64.94 85.94 96.14 86.78 86.94 97.48 84.43 95.02 96.60 79.78
PatchTST [24] 86.83 89.30 66.36 82.75 95.68 87.12 87.19 97.39 83.85 95.00 96.09 83.11
ST-MEM [19] 88.56 92.69 72.88 86.73 96.47 85.17 87.45 97.71 89.70 95.02 95.72 84.51
@@
Supervised [14] - - - 99.26 99.66 99.46 - - 0.7443 0.1481
MAE [25] 99.15 99.56 99.36 99.30 99.62 99.46 0.7398 0.1498 0.7673 0.1089
MTAE [15] 97.62 99.19 98.40 99.17 99.35 99.26 0.7426 0.1100 0.7581 0.1394
MLAE [15] 94.45 98.54 96.45 98.01 98.92 98.46 0.7612 0.1155 0.7912 0.1641
PatchTST [24] 89.91 98.13 98.58 99.12 99.40 99.28 0.7347 0.1089 0.7560 0.1487
ST-MEM [19] 89.93 98.02 97.00 99.19 99.57 99.38 0.7527 0.1230 0.7623 0.0915
@@
Unlike existing SSL methods that primarily focus on classification, we provide a more comprehensive downstream evaluation by additionally conducting segmentation and forecasting tasks on CPSC-QRS [6] and SAMI-TROP [29], respectively. As shown in Table 3, our method achieves substantial improvements across both tasks. Notably, despite the near-saturation performance of baseline methods on the segmentation task, our approach yields consistent and measurable improvements. These improvements stem from our reconstruction of multi-granularity representations in the frequency domain, guiding the model to capture comprehensive diagnostic features and enabling more generalizable feature learning across diverse downstream tasks.
@@ 5.3 Ablation Studies
Having established the presence of simplicity bias in ECG analyses and shown that standard SSL methods provide only partial mitigation, we further conduct a visual comparison between our proposed LEAST and existing ECG SSL approaches. Specifically, we employ CAM to examine how different models attend to critical waveform regions. Using representative samples from CPSC2018 (with lead II as an example), we highlight challenging cases such as AFiB. As illustrated in Fig. 6, we compare MAE [15], ST-MEM [19], and our LEAST. Notably, LEAST demonstrates markedly improved attention to the structured P–QRS–T waveform, effectively identifying the presence of P waves and subtle deviations of F waves from the counterparts. This visual evidence further validates LEAST’s capability to alleviate simplicity bias and enhance fine-grained feature extraction in complex ECG signals.
@@
Prototypical features are known for enhancing feature robustness [30, 31]. In MGPR, feature alignment between encoder and decoder blocks is achieved through prototype representations, providing more stable features compared to batch-level matching. As shown in Table 5, the results demonstrate the superiority of prototype-based matching over batch-level methods. Importantly, the alignment follows a shallow-to-deep strategy based on relative block depth, with selected encoder blocks distributed across varying depths to enhance diversity and avoid concentration. To explore different encoder block configurations, we conduct experiments using our model with 12 encoder blocks and 8 decoder blocks, as shown in Table 6. The results indicate that as long as the shallow-to-deep order is maintained and the selections are well-distributed, variations in block choices have minimal impact on performance. This confirms that MGPR’s alignment strategy is robust, flexible, and effective without rigid block assignments.
@@ Appendix C ECG-adapted Patch Embedding
To overcome this dilemma, we adopt a hybrid approach that leverages the strength of convolutional networks for local pattern extraction and cross-channel interaction, followed by Transformer-based global sequence modeling. Specifically, we employ a lightweight 1D ResNet [9] with convolutional kernel size 5 as a patch embedding module, which processes the input ECG signal of shape $B\times 12\times 1000$ ($batchsize\times leads\times time$). The ResNet encoder reduces the signal to an output of shape $B\times 256\times 125$, where $125$ represents the number of patches (i.e., temporal tokens) and $256$ denotes the embedding dimension for each patch. This architecture allows the model to preserve inter-lead dependencies and extract hierarchical local features, while producing a manageable number of informative patches suitable for Transformer-based attention modeling.

### v4

@@ Abstract
The diagnostic value of electrocardiogram (ECG) lies in its dynamic characteristics, ranging from rhythm fluctuations to subtle waveform deformations that evolve across time and frequency domains. However, supervised ECG models tend to overfit dominant and repetitive patterns, overlooking fine-grained but clinically critical cues—a phenomenon known as Simplicity Bias (SB), where models favor easily learnable signals over subtle but informative ones. In this work, we first empirically demonstrate the presence of SB in ECG analyses and its negative impact on diagnostic performance, while simultaneously discovering that self-supervised learning (SSL) can alleviate it, providing a promising direction for tackling the bias. Following the SSL paradigm, we propose a novel method comprising two key components: 1) Temporal-Frequency aware Filters to capture temporal-frequency features reflecting the dynamic characteristics of ECG signals, and 2) building on this, Multi-Grained Prototype Reconstruction for coarse and fine representation learning across dual domains, further mitigating SB. To advance SSL in ECG analyses, we curate a large-scale multi-site ECG dataset with 1.53 million recordings from over 300 clinical centers. Experiments on three downstream tasks across six ECG datasets demonstrate that our method effectively reduces SB and achieves state-of-the-art performance.
@@ 1 Introduction
The electrocardiogram (ECG) remains a cornerstone of non-invasive cardiac assessment, offering a cost-effective and widely accessible method for monitoring life-threatening conditions such as acute myocardial ischemia bhatia2018screening ; liu2023artificial ; lai2023practical ; yagi2024routine . The beauty of ECG lies in its structured patterns encoded across both temporal and frequency domains. As illustrated in Fig. 1, ECG signals exhibit prominent, stable patterns like QRS complexes and rhythm regularities, characterized as low-frequency features representing dominant coarse-grained characteristics. In contrast, rapidly varying and delicate patterns, such as P and T waves—crucial for detecting early-stage or complex cardiac abnormalities—are identified as high-frequency features, encoding subtle fine-grained diagnostic information. Despite the success of powerful supervised models in advancing ECG analyses wang2024optimizing ; gao2023ecg ; strodthoff2020deep ; alday2020classification ; wang2017time , they often overfit to low-frequency patterns while underrepresenting high-frequency cues critical for clinical interpretation, a phenomenon known as simplicity bias shah2020pitfalls ; teney2022evading in supervised learning paradigm.
Simplicity bias (SB) refers to the tendency of supervised neural networks to prioritize dominant, easily captured features while overlooking subtle but clinically important deviations shah2020pitfalls ; wei2025delving . Conversely, self-supervised learning (SSL) can mitigate SB by encouraging models to learn more comprehensive representations wei2025delving ; sbssl . To investigate whether SB manifests in ECG analyses and whether existing ECG SSL methods can alleviate it, we conduct preliminary experiments. As illustrated in Fig. 2(a), supervised models such as ResNet wang2017time and ViT vaid2023foundational tend to focus primarily on dominant QRS complexes and global pseudo-periodic rhythms, while often neglecting subtle high-frequency components such as P and fibrillatory (F) waves. In contrast, the self-supervised model (MAE zhang2022maefe ) demonstrates partially improved attention to these clinically significant regions, as further discussed in Section 2. Nonetheless, current ECG SSL methods vaid2023foundational ; zhang2022maefe ; hu2023spatiotemporal ; wang2023unsupervised ; lai2023practical ; wang2023adversarial ; na2024guiding are not explicitly designed to address SB. In particular, they lack mechanisms to capture multi-granular representations and model the hierarchical dependencies across temporal-frequency domains, limiting their ability to fully capture the diagnostic richness embedded in ECG signals.
@@
To address such limitation, we propose a novel self-supervised learning method, muLti-grained tEmporal-frequency Aware maSked auToencoders (LEAST). Our proposed method, LEAST, integrates temporal-frequency attention and multi-grained modeling into a masked autoencoder framework, encouraging the self-supervised model to capture challenging features and mitigate SB. The Method consists of two key modules: (1) Temporal-Frequency aware Filters (TFF): We propose multi-head filters within the encoder to capture inherent patterns across both temporal and frequency domains. This design seamlessly integrates temporal and frequency information while capturing domain-specific patterns, enhancing the model’s sensitivity to the dynamic nature of ECG signals and effectively mitigating SB. Building on this, (2) Multi-Grained Prototype Reconstruction (MGPR): To further extend the model’s sensitivity from coarse-grained to fine-grained features, we construct prototypical representations at multiple semantic granularities across dual domains, encouraging the model to preserve information beyond dominant patterns. To advance self-supervised learning research in ECG analyses, we introduce a Chinese multi-site ECG arrhythmia dataset with approximately 380,000 samples collected from 356 clinical centers. Additionally, we integrate multiple publicly available datasets gow2023mimic ; ribeiro2021code to build a large-scale multi-site ECG dataset, comprising a total of 1,530,000 samples.
@@ 2.1 Preliminary Experiments
We perform standard supervised learning on CPSC2018 liu2018open for ECG classification, using ResNet wang2017time as the backbone. Guided by class activation mapping (CAM) cam , we visualize the model’s attention on an atrial fibrillation (AFiB) signal (lead II used) in Fig. 2(a). The results reveal that the supervised model predominantly focuses on easily recognizable QRS complexes, while largely ignoring subtle regions like P waves and F waves, providing clear evidence of SB in ECG analyses.
@@ 2.2 Incorporating Self-Supervised Learning to Mitigate Simplicity Bias
Based on the previous verification of SB in supervised ECG models, we further evaluate whether SSL can alleviate this issue. We adopt a SSL framework in MAE fashion zhang2022maefe and compare it with a supervised baseline, ViT vaid2023foundational , under identical settings (see Appendix A.1). As shown in Fig. 2(a), both models focus on dominant QRS complexes. However, the MAE shows moderately enhanced attention to the presence of P waves, partially mitigating SB. Nonetheless, the improvement remains limited, the ECG SSL model still underattends fine-grained structures critical for diagnosing complex arrhythmias.
@@ 4.2 Complementary Public ECG Datasets
MIMIC-IV-ECG gow2023mimic includes approximately 800,000 12-lead recordings from 160,000 patients collected between 2008–2019 at the Beth Israel Deaconess Medical Center in Boston, USA. As a representative of high-acuity inpatient and ICU care in a developed country, it offers valuable insight into ECG patterns under intensive monitoring.
CODE-15% ribeiro2021code contains 345,779 12-lead ECGs from 233,770 patients across hundreds of municipalities in Brazil. This dataset reflects low-resource outpatient and primary care settings, contributing substantial geographic and epidemiological diversity.
@@ 5.1 Datasets and Empirical Settings
To advance research on self-supervised pretraining in ECG analyses, we constructed large-scale multi-site ECG pretraining dataset. It includes 380,000 proprietary samples collected from over 356 hospitals, fully annotated with 61 diagnostic categories, as well as curated public datasets gow2023mimic ; ribeiro2021code , resulting in a combined corpus of 1,530,000 samples. For fair comparison, all self-supervised baselines zhang2022maefe ; Yuqietal-2023-PatchTST ; mae ; chen2021empirical ; wang2023unsupervised ; na2024guiding are re-pretrained on this dataset in our experiments.
For downstream adaptation, we adopt two paradigms: linear probing and full fine-tuning. We conduct experiments across three downstream tasks—classification, segmentation (detecting characteristic waveform components, i.e. R-peaks locating task) and forecasting (predicting survival time from ECG signals, a form of survival analyses)—on six widely used ECG datasets. More details about the task definition, downstream datasets and metrics could be seen in Appendix B.1. For a fair comparison, following vaid2023foundational ; zhang2022maefe ; na2024guiding , all experiments are conducted using the ViT-Base encoder vaid2023foundational , pretrained on our curated ECG dataset. The empirical settings for training are detailed in the Appendix B.4.
@@ 5.2 Main Results
We conduct classification experiments on PTB-XL strodthoff2020deep , Ningbo zheng2020optimal , Chapman zheng202012 , and CPSC2018 liu2018open , covering a wide range of clinical conditions, geographic regions, and patient demographics. We report the results in Table 2. Results show that our LEAST method consistently achieves significant accuracy improvements over state-of-the-art methods.
@@
MAE mae 87.64 91.37 70.48 73.16 89.25 76.54 73.31 89.36 75.35 94.81 95.68 76.11
MOCO v3 chen2021empirical 86.76 90.92 67.17 73.40 90.12 76.11 74.41 86.13 74.93 92.37 93.70 70.42
MTAE zhang2022maefe 86.79 89.46 66.41 72.89 88.49 76.63 75.33 89.44 76.24 95.15 95.46 73.00
MLAE zhang2022maefe 86.68 88.87 63.88 71.61 88.35 74.35 69.69 88.26 71.12 94.16 94.90 65.67
PatchTST Yuqietal-2023-PatchTST 85.37 86.27 64.54 72.20 87.18 73.41 67.92 85.03 68.87 94.33 95.00 70.38
ST-MEM na2024guiding 87.33 90.30 66.39 73.25 89.22 76.03 73.12 89.91 75.44 95.24 95.42 75.17
@@
MAE mae 88.81 93.00 73.36 86.01 96.64 86.70 88.20 97.33 89.61 96.60 97.23 84.50
MOCO v3 chen2021empirical 87.31 92.27 71.45 86.55 96.42 84.39 85.25 96.61 87.12 95.28 95.29 79.24
MTAE zhang2022maefe 87.39 90.33 69.36 86.31 95.52 87.45 85.94 97.70 85.06 95.37 95.73 83.04
MLAE zhang2022maefe 88.12 90.28 64.94 85.94 96.14 86.78 86.94 97.48 84.43 95.02 96.60 79.78
PatchTST Yuqietal-2023-PatchTST 86.83 89.30 66.36 82.75 95.68 87.12 87.19 97.39 83.85 95.00 96.09 83.11
ST-MEM na2024guiding 88.56 92.69 72.88 86.73 96.47 85.17 87.45 97.71 89.70 95.02 95.72 84.51
@@
Supervised vaid2023foundational - - - 99.26 99.66 99.46 - - 0.7443 0.1481
MAE mae 99.15 99.56 99.36 99.30 99.62 99.46 0.7398 0.1498 0.7673 0.1089
MTAE zhang2022maefe 97.62 99.19 98.40 99.17 99.35 99.26 0.7426 0.1100 0.7581 0.1394
MLAE zhang2022maefe 94.45 98.54 96.45 98.01 98.92 98.46 0.7612 0.1155 0.7912 0.1641
PatchTST Yuqietal-2023-PatchTST 89.91 98.13 98.58 99.12 99.40 99.28 0.7347 0.1089 0.7560 0.1487
ST-MEM na2024guiding 89.93 98.02 97.00 99.19 99.57 99.38 0.7527 0.1230 0.7623 0.0915
@@
Unlike existing SSL methods that primarily focus on classification, we provide a more comprehensive downstream evaluation by additionally conducting segmentation and forecasting tasks on CPSC-QRS gao2023ecg and SAMI-TROP ribeiro2021sami , respectively. As shown in Table 3, our method achieves substantial improvements across both tasks. Notably, despite the near-saturation performance of baseline methods on the segmentation task, our approach yields consistent and measurable improvements. These improvements stem from our reconstruction of multi-granularity representations in the frequency domain, guiding the model to capture comprehensive diagnostic features and enabling more generalizable feature learning across diverse downstream tasks.
@@ 5.3 Ablation Studies
Having established the presence of simplicity bias in ECG analyses and shown that standard SSL methods provide only partial mitigation, we further conduct a visual comparison between our proposed LEAST and existing ECG SSL approaches. Specifically, we employ CAM to examine how different models attend to critical waveform regions. Using representative samples from CPSC2018 (with lead II as an example), we highlight challenging cases such as AFiB. As illustrated in Fig. 6, we compare MAE zhang2022maefe , ST-MEM na2024guiding , and our LEAST. Notably, LEAST demonstrates markedly improved attention to the structured P–QRS–T waveform, effectively identifying the presence of P waves and subtle deviations of F waves from the counterparts. This visual evidence further validates LEAST’s capability to alleviate simplicity bias and enhance fine-grained feature extraction in complex ECG signals.
@@
Prototypical features are known for enhancing feature robustness snell2017prototypical ; duan2024long . In MGPR, feature alignment between encoder and decoder blocks is achieved through prototype representations, providing more stable features compared to batch-level matching. As shown in Table 5, the results demonstrate the superiority of prototype-based matching over batch-level methods. Importantly, the alignment follows a shallow-to-deep strategy based on relative block depth, with selected encoder blocks distributed across varying depths to enhance diversity and avoid concentration. To explore different encoder block configurations, we conduct experiments using our model with 12 encoder blocks and 8 decoder blocks, as shown in Table 6. The results indicate that as long as the shallow-to-deep order is maintained and the selections are well-distributed, variations in block choices have minimal impact on performance. This confirms that MGPR’s alignment strategy is robust, flexible, and effective without rigid block assignments.
@@ Appendix C ECG-adapted Patch Embedding
To overcome this dilemma, we adopt a hybrid approach that leverages the strength of convolutional networks for local pattern extraction and cross-channel interaction, followed by Transformer-based global sequence modeling. Specifically, we employ a lightweight 1D ResNet [9] with convolutional kernel size 5 as a patch embedding module, which processes the input ECG signal of shape $B\times 12\times 1000$ ($batchsize\times leads\times time$). The ResNet encoder reduces the signal to an output of shape $B\times 256\times 125$, where 125125125 represents the number of patches (i.e., temporal tokens) and 256256256 denotes the embedding dimension for each patch. This architecture allows the model to preserve inter-lead dependencies and extract hierarchical local features, while producing a manageable number of informative patches suitable for Transformer-based attention modeling.